### v1

# Colorful circuits and colorful topes in oriented matroids

Abstract.
We provide a short proof of a conic version of the colorful Carathéodory theorem for oriented matroids. Holmsen’s extension of the colorful Carathéodory theorem to oriented matroids (Advances in Mathematics, 2016) already encompasses several generalizations of the original result, but not its conic version. Our approach relies on a common generalization of Sperner’s lemma and Meshulam’s lemma—two closely related results from combinatorial topology that have found a number of applications in discrete geometry and combinatorics. This generalization may be of independent interest.
Using a similar approach, we also establish the following colorful theorem for topes, whose special geometric case had not been considered before: Given $n$ topes from a uniform oriented matroid with $n$ elements, if they agree on some element, then there is a way to select a distinct element from each tope, together with its sign, so as to form another tope of the oriented matroid. Motivated by this theorem, we further explore other conditions leading to the same conclusion.
Key words and phrases:
Oriented matroids; colorful Carathéodory theorem; Sperner lemma; topes
2020 Mathematics Subject Classification:
05B35, 52C40
The first author was supported by a KIAS Individual Grant (CG102101) at Korea Institute for Advanced Study.
The second author was supported by the Institute for Basic Science (IBS-R029-C1) and the Department of Mathematical Sciences of Korea Advanced Institute of Science and Technology (BK21).
1. Introduction
One of the most beautiful theorems in discrete geometry is the colorful Carathéodory theorem by Bárány [2]. It states: Given $d+1$ sets of points in $\mathbb{R}^{d}$, each of them containing the origin in its convex hull, there exists a way to pick one point from each set so that the resulting set still contains the origin in its convex hull. A generalization for oriented matroids has been proved by Holmsen [7] via a generalization of the topological colorful Helly theorem by Kalai and Meshulam [8]. Actually, the result by Holmsen (as well as that by Kalai and Meshulam) is a generalization of a much stronger version of the colorful Carathéodory theorem, where the colors are replaced by some (non-oriented) matroid constraint, and where the condition is less demanding. Yet, there exists a “conic” version of the colorful Carathéodory theorem: Given $d+1$ sets of vectors in $\mathbb{R}^{d+1}$, and a vector $p$ in $\mathbb{R}^{d+1}$ contained in the conic hull of each of these sets, there exists a way to pick one vector from each set so that the resulting set still contains $p$ in its conic hull. This conic version is also stronger than the original convex version.
A first contribution of this paper is a generalization of the conic version for oriented matroids, something which was absent from the literature. Its proof—given in Section 2.3—relies on a topological lemma, which turns out to be a simultaneous generalization of the Sperner lemma [14] and the Meshulam lemma [10], both lemmas which have found many applications in topological combinatorics, especially to get “colorful” combinatorial results (e.g., [1, 11]), and beyond [5]. Such a generalization is new, and may be of independent interest. It is stated and proved in Section 2.1.
Theorem 1.1.
Let $C_{1},\dots,C_{r}$ be positive circuits of an oriented matroid of rank $r$ that all contain a common element $e$. Then there exist elements $f_{i}\in C_{i}\setminus\{e\}$  for $i=1,\dots,r$ such that the set $\{e,f_{1},\dots,f_{r}\}$ contains a positive circuit with $e$ as element.
A generalization of the convex colorful Carathéodory theorem for oriented matroids can easily be obtained from Theorem 1.1 by a lifting argument. For sake of completeness, we state and prove this generalization. (Note that it is actually a special case of Holmsen’s theorem.)
Theorem 1.2.
Let $C_{1},\dots,C_{r}$ be positive circuits of an oriented matroid of rank r1r-1. Then there exist elements $f_{i}\in C_{i}$  for $i=1,\dots,r$ such that the set $\{f_{1},\dots,f_{r}\}$ contains a positive circuit.
Proof (based on Theorem 1.1).
Let $\sigma$ be a localization of the dual oriented matroid. We choose $\sigma$ in such a way that $\sigma(C)=+$ for every positive circuit of the original oriented matroid. This is possible thank to a result by Las Vergnas [4, Proposition 7.2.4] (applied with all elements of the ground set; this result requires to specify the signs taken by the localization on those elements via a vector $\alpha$,” which we set to the all-++ vector). Apply then Theorem 1.1 to the circuits $(C_{i},\sigma(C_{i}))$, with $e$ being the new element. This leads to elements $f_{i}\in C_{i}$ such that $\{e,f_{1},\ldots,f_{r}\}$ contains a positive circuit of the lifted oriented matroid. Contracting $e$ shows that $\{f_{1},\ldots,f_{r}\}$ contains a positive circuit of the original oriented matroid. ∎
A second contribution is a result with a similar flavor. Yet, as far as we know, even the special case when the oriented matroid is realizable is new. Its proof—given in Section 2.4—also relies on the Sperner–Meshulam-type lemma of Section 2.1.
Note that we are moving to the dual oriented matroid (topes are covectors while circuits are vectors). We chose to state (and prove) the theorems in this way to keep the statements as concise and natural as possible.
Theorem 1.3.
Consider a uniform oriented matroid on $[n]$. For every collection of $n$ topes $T_{1},T_{2},\ldots,T_{n}$ that agree on element $n$, there exists a tope $\widetilde{T}$ agreeing with each $T_{i}$ on a distinct element.
While Theorems 1.1 and 1.2 hold without any assumption on non-degeneracy, uniformity is required for Theorem 1.3, as shown by the following example. Consider the covectors of the matrix $\begin{pmatrix}1&1&1&0\\ 0&0&0&1\end{pmatrix}$. They form a rank-two oriented matroid on four elements, with
[…]
as topes. Though, there is no $\widetilde{T}$ as in the theorem for $T_{1}=T_{2}=(+,+,+,+)$ and $T_{3}=T_{4}=(-,-,-,+)$. (This example can immediately be generalized to get a counter-example of any rank.)
It is not clear that the condition in Theorem 1.3 about element $n$ is necessary, but just removing this condition is not possible. Indeed, the (uniform) oriented matroid with $n$ elements and with only two cocircuits, (+,+,,+)($\ldots$,+) and $(-,-,\ldots,-)$, is a counter-example. However, finding other counter-examples does not look that easy and this leads us to formulate the following open question.
Open Question 1.
Given $n$ topes $T_{1},\dots,T_{n}$ of a uniform oriented matroid on $n$ elements, under what conditions does there exist a tope $\widetilde{T}$ agreeing with each tope $T_{i}$ on a distinct element?
We did not succeed in answering this question in full generality. However, we were able to identify two sufficient conditions, in addition to that of Theorem 1.3, ensuring a positive answer to Question 1. One condition is that, among the $n$ topes, at most $r$ are distinct, where $r$ is the rank of the oriented matroid. It is stated as follows and the proof is given in Section 2.4 as well.
Theorem 1.4.
Consider a uniform oriented matroid of rank $r$ on $[n]$. For every collection of topes $T_{1},T_{2},\ldots,T_{r}$ and positive integers $n_{1},n_{2},\ldots,n_{r}$ summing up to $n$, there exist a tope $\widetilde{T}$ and a partition $n]=J_{1}\cup J_{2}\cup\cdots\cup J_{r}$ such that $|J_{i}|=n_{i}$ and $\widetilde{T}(j)=T_{i}(j)$ for all $i\in[r]$ and $j\in J_{i}$.
The rank of the oriented matroid being two is the second sufficient condition we have been able to identify. The proof—given in Section 3—relies on “alternating words” on the alphabet $\{+,-\}$.
Theorem 1.5.
Consider a uniform oriented matroid of rank two on $n$ elements. For every collection of $n$ topes $T_{1},T_{2},\ldots,T_{n}$, there exists a tope $\widetilde{T}$ agreeing with each $T_{i}$ on a distinct element.
Notice that the counter-example above, showing that extra conditions are indeed necessary to get a positive answer to Question 1, is of rank one. So, in the statement of Theorem 1.5, it is not possible to replace “of rank two” by “of rank at most two.”
Conventions
[…]
Acknowledgments
The authors thank Eran Nevo for helping them write the argument relying on Alexander’s duality in the proof of Lemma 2.3. They are also grateful to Laura Anderson for fruitful discussions about triangulations of pseudosphere representations and fundamental facts of oriented matroids, and to Andreas Holmsen for explanations regarding the status of the colorful Carathéodory theorem within oriented matroids. Last but not least, they thank Michael Gene Dobbins for helpful discussion and comments regarding Open Question 1.
2. Proofs of Theorems 1.11.3, and 1.4
The proofs of Theorems 1.11.3, and 1.4 share similar ingredients. They all use the Sperner–Meshulam-type lemma mentioned in the introduction. To apply this lemma, connectivity results for specific simplicial complexes are also required. Such results are obtained with the help of the topological representation theorem.
2.1. A “Sperner–Meshulam” lemma
Lemma 2.1.
Consider a positive integer $k$ and a family $\{\mathsf{K}^{I}\}_{I\subseteq[k]}$ of simplicial complexes such that $\mathsf{K}^{I}$ is a subcomplex of $\mathsf{K}^{J}$ whenever $I\subseteq J$. Suppose that $\mathsf{K}^{I}$ is homologically (|I|2)(|I|-2)-connected for every $I$, and assume we are given a labeling $\lambda$ of the vertices of $\mathsf{K}^{[k]}$ such that $\lambda(v)\in I$ whenever $v\in\mathsf{K}^{I}$. Then there exists a (k1)(k-1)-dimensional simplex of $\mathsf{K}^{[k]}$ whose vertices get distinct labels by $\lambda$.
The Sperner lemma is the special case when $\mathsf{K}^{[k]}$ is a triangulation of a standard simplex with $[k]$ as vertex set and $\mathsf{K}^{I}$ is the triangulation induced on the face of the standard simplex with $I$ as vertex set. The Meshulam lemma is the special case when $\mathsf{K}^{I}$ is the subcomplex of $\mathsf{K}^{[k]}$ induced by the vertices labeled with elements from $I$.
Proof of Lemma 2.1.
We start by showing the existence of a chain $c^{I}\in C_{|I|-1}(\mathsf{K}^{[k]},\mathbb{Z}_{2})$ for each subset $I$ of $[k]$ such that $c^{\varnothing}=1$ and, for every non-empty subset $I$ of $[k]$, the following two properties hold:
  1. (i)
    $\partial c^{I}=\sum_{i\in I}c^{I\setminus\{i\}}$.
  2. (ii)
    $c^{I}$ is supported by $\mathsf{K}^{I}$.
(When $|I|=1$, we see $\partial$ as the standard augmentation map used to define reduced homology.) To do so, we set $c^{\varnothing}\coloneqq 1$ and $c^{\{i\}}\coloneqq v_{i}$ where $v_{i}$ is an arbitrary vertex of $\mathsf{K}^{\{i\}}$. (Note that such a vertex exists since being (1)(-1)-connected means being non-empty.) The relation $\partial c^{\{i\}}=c^{\varnothing}$ holds. Suppose now that for all $I$ of size at most $s\leqslant k-1$ the chains $c^{I}$ have been defined and that they satisfy the two properties (i) and (ii). Consider now $I^{\prime}\subseteq[k]$ of size $s+1$. We have
$\partial\sum_{i\in I^{\prime}}c^{I^{\prime}\setminus\{i\}}=\sum_{i\in I^{\prime}}\partial c^{I^{\prime}\setminus\{i\}}=\sum_{i\in I^{\prime}}\sum_{i^{\prime}\in I^{\prime}\setminus\{i\}}c^{I^{\prime}\setminus\{i,i^{\prime}\}}=0\,$,
where induction is used to get the second equality. The simplicial complex $\mathsf{K}^{I^{\prime}}$ is homologically $(|I^{\prime}|-2)$-connected, and thus there exists a chain $c^{I^{\prime}}$ supported by $\mathsf{K}^{I^{\prime}}$, which implies (ii), such that $\partial c^{I^{\prime}}=\sum_{i\in I^{\prime}}c^{I^{\prime}\setminus\{i\}}$, which implies (i).
Then, we interpret the labeling $\lambda$ as a simplicial map from $\mathsf{K}^{[k]}$ to the standard simplex with $[k]$ as vertex set, which we denote by $\sigma^{[k]}$. Since $\lambda(v)\in I$ when $v\in V(\mathsf{K}^{I})$ and since the $c^{I}$ are supported by $\mathsf{K}^{I}$, there exist $\alpha_{I}\in\mathbb{Z}_{2}$ such that $\lambda_{\sharp}(c^{I})=\alpha_{I}\sigma^{I}$ for all $I\subseteq[k]$ (where $\sigma^{I}$ is the standard simplex with $I$ as vertex set). Note that $\alpha_{\{i\}}=1$ for all $i\in[k]$. By the definition of the $c^{I}$ and the property of chain maps, we have $\partial\lambda_{\sharp}(c^{I})=\sum_{i\in I}\lambda_{\sharp}(c^{I\setminus\{i\}})$. Since $\partial\sigma^{I}=\sum_{i\in I}\sigma^{I\setminus\{i\}}$, a direct induction shows that $\alpha_{I}=1$ for every non-empty subset $I$ of $[k]$. There exists thus a (k1)(k-1)-dimensional simplex of $\mathsf{K}^{[k]}$ whose vertices get distinct labels by $\lambda$. This translates into the desired conclusion. ∎
2.2. Preliminary results
In this subsection, we consider an oriented matroid $\mathcal{M}$ on $[n]$. We establish and present elementary results on $\mathcal{V}$ and $\mathcal{L}$—its sets of vectors and covectors—seen as partially ordered sets (with the usual order of oriented matroids). The proofs will rely on the topological representation theorem; see [4, Chapter 5]. We emphasize that because of the choice we made regarding the statements of the theorems—with circuits for Theorem 1.1 and with topes for Theorems 1.3 and 1.4—the topological representation theorem will be used sometimes for $\mathcal{M}$ and sometimes for its dual.
The set of vectors of an oriented matroid forms a lattice. We denote by $\operatorname{height}(X)$ the height of a vector $X$, defined as the size of a maximal chain from the zero vector to $X$, minus one. In particular, the height of the zero vector is zero and the height of any circuit is one.
Proposition 2.2.
The inequality $|\underline{X}|\leqslant\operatorname{rk}(\mathcal{M})+\operatorname{height}(X)$ holds for every vector $X$.
Proof.
Let $\mathcal{M}^{\prime}$ be the oriented matroid obtained from $\mathcal{M}$ by contracting all coloops. Notice that the rank decreases when coloops are contracted, but that vectors are not modified. Consider an essential topological representation of the dual of $\mathcal{M}^{\prime}$. The dimension of the cell associated with a vector $X$ has dimension $\operatorname{height}(X)-1$. The dimension of the representation being $n-\operatorname{rk}(\mathcal{M}^{\prime})-1$, the cell $X$ is contained in at least $(n-\operatorname{rk}(\mathcal{M})-1)-(\operatorname{height}(X)-1)=n-\operatorname{rk}(\mathcal{M}^{\prime})-\operatorname{height}(X)$ pseudospheres, i.e., the number of zeroes in $X$ is at least $n-\operatorname{rk}(\mathcal{M}^{\prime})-\operatorname{height}(X)$. The size of $\underline{X}$ is at most $\operatorname{rk}(\mathcal{M}^{\prime})+\operatorname{height}(X)$. Since $\operatorname{rk}(\mathcal{M}^{\prime})\leqslant\operatorname{rk}(\mathcal{M})$, we get the desired inequality. ∎
We are going to state and prove a crucial lemma for the proof of Theorem 1.1. For an element $e$ of $\mathcal{M}$ and a positive integer $h$, define $\mathcal{V}^{+}_{e,h}$ as the set of positive vectors $X$ satisfying simultaneously the following two conditions:
  • $C\preccurlyeq X\,\Longrightarrow\,C(e)=+$ for every circuit $C$.
  • $1\leqslant\operatorname{height}(X)\leqslant h$.
Lemma 2.3.
If $\mathcal{V}^{+}_{e,h}$ is not empty, then its order complex is homologically (h2)(h-2)-connected.
Proof.
We assume that $\mathcal{V}^{+}_{e,h}$ is not empty. In particular, $e$ is not a coloop. Since no vector contains a coloop, we can actually assume that $\mathcal{M}$ has no coloop. Take an essential topological representation of the dual of $\mathcal{M}$, where we denote as usual by $S_{j}$ the pseudosphere associated with an element $j$. In this proof, we work with the cell complex $X$ induced by this topological representation. By default, we consider closed cells.
Consider the cell $\sigma$ corresponding to the composition of all positive circuits. Denote by $U^{+}$ the vertices of $\sigma$ that are not on $S_{e}$ (which is not empty because we assume that $\mathcal{V}^{+}_{e,h}$ is not empty), and by $U^{0}$ the vertices that are on $S_{e}$. There are two possibilities. For each of them, we will consider the barycentric subdivision of a subcomplex of $X$. Such a subdivision exists because it is a regular cell complex [4, Proposition 4.7.8].
The first possibility is that $U^{0}=\varnothing$. In this case, the order complex of $\mathcal{V}^{+}_{e,h}$ is the barycentric subdivision of the (h1)(h-1)-skeleton of $\sigma$, and thus homologically (h2)(h-2)-connected.
The second possibility is that $U^{0}\neq\varnothing$. Set $h^{\prime}\coloneqq\min(h,\dim(\sigma))$. The order complex of $\mathcal{V}^{+}_{e,h}$ is the barycentric subdivision of the $(h^{\prime}-1)$-skeleton of $(\partial\sigma)[U^{+}]$ (subcomplex of $\partial\sigma$ induced by $U^{+}$). We finish the proof by establishing that $(\partial\sigma)[U^{+}]$ is acyclic: by the definition of homology for cell complexes, this implies immediately that the barycentric subdivision of its $(h^{\prime}-1)$-skeleton is homologically $(h^{\prime}-2)$-connected, and acyclic when $h>\dim(\sigma)$.
We take the barycentric subdivision of $\partial\sigma$, whose existence is guaranteed by $X$ being a regular cell complex. We denote by $\operatorname{sd}(\partial\sigma)$ this barycentric subdivision. Consider the collection of stars $\operatorname{st}(u)$ for $u$ ranging $U^{0}$. Here, we consider “closed stars”, i.e., the collection of simplices $\tau$ in $\operatorname{sd}(\partial\sigma$) that together with $u$ form a simplex of $\operatorname{sd}(\partial\sigma)$. We make now two crucial remarks. The first one is that these stars are simplicial subcomplexes. The second one goes as follows: for every subset $U^{\prime}\subseteq U^{0}$, either there is no cell of $\partial\sigma$ containing $U^{\prime}$ completely, in which case $\bigcap_{u\in U^{\prime}}\operatorname{st}(u)=\varnothing$, or there is a (unique) minimal cell $\tau$ containing $U^{\prime}$, in which case $\bigcap_{u\in U^{\prime}}\operatorname{st}(u)$ has all its maximal simplices containing the barycenter of $\tau$. (Note that we use here the lattice structure of the face poset of $\sigma$.) In other words, if $\bigcap_{u\in U^{\prime}}\operatorname{st}(u)$ is non-empty, it is a cone with apex the barycenter of $\tau$, and thus contractible. These two remarks make that we can apply the nerve theorem, in a version for simplicial subcomplexes [3, Theorem 6]: homology of $\bigcup_{u\in U^{0}}\operatorname{st}(u)$ is that of the nerve of the stars $\operatorname{st}(u)$ for $u$ ranging over $U^{0}$. Since $\bigcap_{u\in U^{0}}\operatorname{st}(u)$ contains the barycenter of the cell $\sigma\cap S_{e}$, it is non-empty, which means that the nerve of the stars is a full simplex, and this implies that $\bigcup_{u\in U^{0}}\operatorname{st}(u)$ is acyclic. We finish by the following chain of equalities and homotopy equivalence. We denote by $B$ the vertex set of $\operatorname{sd}(\partial\sigma)$ and by $B^{0}$ its subset corresponding to barycenters of cells hitting $U^{0}$:
$\|\partial\sigma\|\setminus\bigl{\|}\bigcup_{u\in U^{0}}\operatorname{st}(u)\bigr{\|}=\|\partial\sigma\|\setminus\|\operatorname{sd}(\partial\sigma)[B^{0}]\|\simeq\|\operatorname{sd}(\partial\sigma)[B\setminus B^{0}]\|=\|(\partial\sigma)[U^{+}]\|\,$,
where the two equalities are immediate and where the homotopy equivalence result from one space being a strong deformation retract of the other [4, Lemma 4.7.27]. By Alexander’s duality [12, Theorem 71.1], the acyclicity of $\bigcup_{u\in U^{0}}\operatorname{st}(u)$ implies that of $\|\partial\sigma\|\setminus\bigl{\|}\bigcup_{u\in U^{0}}\operatorname{st}(u)\bigr{\|}$, and thus that of $(\partial\sigma)[U^{+}]$. ∎
The next two lemmas will be used in the proofs of Theorems 1.3 and 1.4. For $J^{+},J^{-}$ two disjoint subsets of $[n]$, set
[…]
Lemma 2.4.
If $J^{+}$ and $J^{-}$ are not both empty and $\mathcal{L}_{J^{+},J^{-}}$ contains a tope, then the order complex of $\mathcal{L}_{J^{+},J^{-}}$ is a ball. If $J^{+}$ and $J^{-}$ are both empty, then the order complex of $\mathcal{L}_{J^{+},J^{-}}$ is an $(\operatorname{rk}(\mathcal{M})-1)$-dimensional sphere.
Proof.
Consider an essential topological representation of $\mathcal{M}$, where we denote as usual the pseudosphere associated with an element $j$ by $S_{j}$, and the two closed hemispheres with $S_{j}$ as boundary respectively $S_{j}^{+}$ and $S_{j}^{-}$. Without loss of generality, we can assume that $\mathcal{M}$ has no loop. The topological representation theorem  ensures the following.
  • When $J^{+}$ and $J^{-}$ are not both empty, the covectors in $\mathcal{L}_{J^{+},J^{-}}$ correspond to the cells situated in $\bigl{(}\bigcap_{j\in J^{+}}S_{j}^{+}\bigr{)}\cap\bigl{(}\bigcap_{j\in J^{-}}S_{j}^{-}\bigr{)}$, which, according to the topological representation theorem, is homeomorphic either to a sphere of dimension at most $\operatorname{rk}(\mathcal{M})-2$, or to a ball. Since it contains a tope, it has to be a ball.
  • When $J^{+}$ and $J^{-}$ are both empty, the covectors in $\mathcal{L}_{J^{+},J^{-}}$ correspond to all cells in $S^{\operatorname{rk}(\mathcal{M})-1}$.∎
The following lemma is a classical result from topological combinatorics. For applications and references to proofs of Quillen’s lemma, see [4, Chapter 4].
Lemma 2.5 (Quillen [13]).
Let $f\colon P\to Q$ be an order-preserving or order-reversing mapping of posets. Suppose that $f^{-1}(Q_{\succcurlyeq x})$ is contractible for all $x\in Q$. Then $f$ induces homotopy equivalence of $\|P\|$ and $\|Q\|$.
2.3. Proof of Theorem 1.1
Proof of Theorem 1.1.
If $e$ is a loop, then the statement is obvious. So assume that $e$ is not a loop. Up to making parallel copies of the elements, we can assume that $e$ is the only element of $\underline{C}_{i}\cap\underline{C}_{j}$ whenever $i\neq j$. Moreover, we can assume that the ground set is $\bigcup_{i=1}^{r}\underline{C}_{i}$. (Note that now $\operatorname{rk}(\mathcal{M})$ is not necessarily $r$, but certainly not bigger.) For $I\subseteq[r]$, let $\mathsf{K}^{I}$ be the order complex of the set of positive vectors $X$ satisfying simultaneously the following three conditions:
  • $C\preccurlyeq X\,\Longrightarrow\,C(e)=+$ for every circuit $C$.
  • $1\leqslant\operatorname{height}(X)\leqslant|I|$.
  • $\underline{X}\subseteq\bigcup_{i\in I}\underline{C}_{i}$.
It is homologically (|I|2)(|I|-2)-connected by Lemma 2.3, applied to the oriented matroid restricted to the ground set $\bigcup_{i\in I}\underline{C}_{i}$, with $h=|I|$. Define a labeling $\lambda$ of the positive vectors of $\mathcal{M}$ by elements from $[r]$ as follows: for a positive vector $X$, let $\lambda(X)$ be the index $i\in[r]$ such that $X$ selects the maximum number of elements of $C_{i}$; in case of a tie, choose the smallest possible $i$. This definition makes that $\lambda(X)\in I$ when $X\in\mathsf{K}^{I}$. (Here, we use that $e$ is not a loop.) The family $\mathsf{K}^{I}$ and the labeling $\lambda$ satisfy the conditions of Lemma 2.1. Hence, there exist positive vectors $X_{1}\prec X_{2}\prec\cdots\prec X_{r}$ such that $\lambda(X_{i})$ get distinct values. By construction, the height of $X_{r}$ is at most $r$, which implies by Proposition 2.2 that $|\underline{X}_{r}|\leqslant 2r$ (Note that here we use $\operatorname{rk}(\mathcal{M})\leqslant r$.) The element $e$ is in the support of all $X_{i}$ by definition of the $\mathsf{K}^{I}$. Thus, there is an $i^{\star}$ such that circuit $C_{i^{\star}}$ contributes in $X_{r}$ by at most one element (apart from $e$). There is a vector $X_{j}$ whose image by $\lambda$ is $i^{\star}$. Necessarily, no circuit $C_{i}$ contributes by more than one to this vector $X_{j}$, and since $X_{1}\preccurlyeq X_{j}$, no circuit $C_{i}$ contributes by more than one to $X_{1}$. This latter vector $X_{1}$ being of height $1$, it is a positive circuit that shares at most one element (outside $e$) with each $C_{i}$, as desired. ∎
2.4. Proof of Theorems 1.3 and 1.4
The proofs of Theorems 1.3 and 1.4 rely on a some specific simplicial complex which we will define now. This simplicial complex depends on a parameter $k$ that will be set to $n$ for the proof of Theorem 1.3 and to $\operatorname{rk}(\mathcal{M})$ for the proof of Theorem 1.4.
Consider a uniform oriented matroid $\mathcal{M}$ on $[n]$ as in the statement of these theorems. Let $T_{1},T_{2},\ldots,T_{k}$ be topes. For a non-empty subset $I$ of $[k]$, we set
$F^{I}\coloneqq\bigl{\{}(i_{1},i_{2},\ldots,i_{n})\in I^{n}\colon(T_{i_{1}}(1),T_{i_{2}}(2),\ldots,T_{i_{n}}(n))\in\mathcal{L}\bigl{\}}\,.$
Let $\sigma^{I}$ be a geometric simplex whose vertices are identified with $I$. Then $(\sigma^{I})^{n}$ is a polytope (a.k.a. simplotope or prodsimplex) whose vertices are all possible $n$-tuples of elements in $I$. Some of these vertices are elements of $F^{I}$. We denote by $\mathsf{L}^{I}$ the polyhedral complex these vertices induce (keeping a face of $(\sigma^{I})^{n}$ only if all its vertices are in $F^{I}$).
We illustrate this construction with $\mathcal{M}$ being the alternating oriented matroid of rank $3$ on $8$ elements (see [4, Section 9.4] for discussions on alternating oriented matroids). For such an oriented matroid, topes correspond to sequences of ++ and - in which the sign changes at most twice. Figure 1 shows an example with $k=3$ topes. The highlighted cells form another tope, showing that $(2,3,3,1,2,3,1,2)$ belongs to $F^{I}$.
[…]
Figure 1. Illustration of the construction of $F^{I}$.
The proof of the theorems relies on the high connectivity of $\mathsf{L}^{I}$, as established by the following lemma. This is the place where the uniformity of $\mathcal{M}$ is needed.
Lemma 2.6.
Suppose that $\mathcal{M}$ is uniform, and let $I$ be a non-empty subset of $[k]$. Then $\mathsf{L}^{I}$ is (r2)(r-2)-connected, where $r$ is the rank of the oriented matroid. Moreover, if there is a $j$ in $[n]$ such that the $T_{i}(j)$ for $i\in I$ are all equal, then $\mathsf{L}^{I}$ is contractible.
Proof.
Let $J^{+}\coloneqq\bigcap_{i\in I}T_{i}^{+}$ and $J^{-}\coloneqq\bigcap_{i\in I}T_{i}^{-}$. Here $T_{i}^{+}$ and $T_{i}^{-}$ denote the sets of elements $j$ such that $T_{i}(j)=+$ and $T_{i}(j)=$-, respectively. We are going to prove that $\mathsf{L}^{I}$ and the order complex $\Delta(\mathcal{L}_{J^{+},J^{-}})$ are homotopy equivalent. By Lemma 2.4, this will be enough. For $\tau=\tau_{1}\times\tau_{2}\times\cdots\times\tau_{n}$ a face of $\mathsf{L}^{I}$, let $X$ be defined by
[…]
Since every vertex of $\tau$ corresponds to a tope, repeated application of vector elimination, in its strong version by Edmonds, Fukuda, and Mandel [6, 9] (we use actually [4, Corollary 3.7.9, Chapter 3]), shows that this $X$ is a covector of $\mathcal{M}$. Set $f(\tau)\coloneqq X$. This defines an order-reversing map $f\colon\Delta(\mathsf{L}^{I})\to\mathcal{L}_{J^{+},J^{-}}$. Given $Y\in\mathcal{L}_{J^{+},J^{-}}$, define $\tau^{Y}=\tau_{1}^{Y}\times\tau_{2}^{Y}\times\cdots\times\tau_{n}^{Y}$ by setting $\tau_{j}^{Y}\coloneqq\sigma^{I}$ if $Y(j)=0$ and by $\tau_{j}^{Y}\coloneqq\sigma^{\{i\in I\colon T_{i}(j)=Y(j)\}}$ otherwise. The oriented matroid $\mathcal{M}$ being uniform, $\tau^{Y}$ is a face of $\mathsf{L}^{I}$. The faces $\tau$ of $\mathsf{L}^{I}$ such that $f(\tau)\succcurlyeq Y$ are precisely the faces of $\tau^{Y}$. Since this latter is a polytope, it is contractible. The desired result follows then from the Quillen lemma (Lemma 2.5). ∎
The proofs of Theorems 1.3 and 1.4 are then short and almost identical.
Proof of Theorem 1.3.
In this proof, we work with $k=n$. Let $\tau=\tau_{1}\times\cdots\times\tau_{n}$ be a face of $\mathsf{L}^{[n]}$ of dimension at most n1n-1. Define $\lambda(\tau)$ as the smallest index $i$ such that there is at least one index $j$ with $\tau_{j}=\{i\}$. Such an $i$ exists because the dimension of $\tau$ is at most n1n-1, which means that there are at least one index $j$ such that $\tau_{j}$ is a singleton. Thanks to Lemma 2.6 and by the definition of $\lambda$, the condition of Lemma 2.1 is satisfied for $\mathsf{K}^{I}$ being the barycentric subdivision of the (|I|1)(|I|-1)-skeleton of $\mathsf{L}^{I}$. This means that there exist simplices $\tau^{1}\subset\cdots\subset\tau^{n}$ of $\mathsf{L}^{[n]}$ such that $\lambda(\tau^{i})=i$ for every $i\in[n]$, with $\tau^{1}$ of dimension 0. The simplex $\tau^{1}$ corresponds thus to a vertex $(i_{1},i_{2},\ldots,i_{n})$, where the $i_{j}$ are formed by the integers $1$ to $n$. ∎
Proof of Theorem 1.4.
In this proof, we work with $k=r=\operatorname{rk}(\mathcal{M})$. Let $\tau=\tau_{1}\times\cdots\times\tau_{n}$ be a face of $\mathsf{L}^{[r]}$ of dimension at most r1r-1. Define $\lambda(\tau)$ as the smallest index $i$ such that there are at least $n_{i}$ indices $j$ with $\tau_{j}=\{i\}$. Such an $i$ exists because the dimension of $\tau$ implies that there are at least $n-r+1>\sum_{i=1}^{r}(n_{i}-1)$ indices $j$ such that $\tau_{j}$ is a singleton. Thanks to Lemma 2.6 and by the definition of $\lambda$, the condition of Lemma 2.1 is satisfied for $\mathsf{K}^{I}$ being the barycentric subdivision of the (|I|1)(|I|-1)-skeleton of $\mathsf{L}^{I}$. This means that there exist simplices $\tau^{1}\subset\cdots\subset\tau^{r}$ of $\mathsf{L}^{[r]}$ such that $\lambda(\tau^{i})=i$ for every $i\in[r]$, with $\tau^{1}$ of dimension 0. Setting $J_{i}$ as the indices $j$ such that $\tau_{j}^{1}=\{i\}$ (where $\tau^{1}=\tau_{1}^{1}\times\cdots\times\tau_{r}^{1}$), we get the desired partition. ∎
3. Alternating words and proof of Theorem 1.5
By a binary word, we mean a word on the alphabet $\{-,+\}$. Such a word is alternating if every two consecutive symbols are opposite. The alternation number of a binary word is the maximal length of an alternating subword (subsequence of non-necessarily consecutive symbols).
Theorem 3.1.
Consider $n$ binary words $w^{1},w^{2},\ldots,w^{n}$ of length $n$. If each $w^{i}$ has alternation number at most two, then there exists a permutation $\pi$ of $[n]$ such that the binary word $w_{1}^{\pi(1)},w_{2}^{\pi(2)},\ldots,w_{n}^{\pi(n)}$ has alternation number at most two.
It is still very possible that versions of Theorem 3.1 for other alternation numbers are true as well (but not for alternation number equal to one). This would contribute to Open question 1. Indeed, binary words of length $n$ with alternation number at most $r$, seen as $n$-tuples, are exactly the topes of the so-called alternating oriented matroid of rank $r$ [4, Section 8.2 and Section 9.4]. When the rank is two, the converse is also true: up to reorientation, every uniform oriented matroid of rank two is alternating. Theorem 1.5 is therefore just a reformulation of Theorem 3.1. (Indeed, up to reorientation, there is essentially a unique topological representation of a uniform oriented matroid of rank two: the representation being on a one-dimensional sphere, it induces a cyclic order on the elements, from which an orientation making it alternating is immediate.)
The remaining of the section is devoted to the proof of Theorem 3.1. The proof relies on two ingredients. The first one is a specific total order on the set of binary words of fixed length and with alternation number at most two. The second one is an elementary topological fact in dimension two.
3.1. A total order on binary words with alternation number two
Consider two words $w,w^{\prime}$ of length $n$ and with alternation number at most two. We set $w\preccurlyeq_{b}w^{\prime}$ if we are in one of the three following (exclusive) situations:
  • $w_{n}=+$ and $w_{n}^{\prime}=$-.
  • $w_{n}=w^{\prime}_{n}=+$ and $(w_{i}=-\Rightarrow w_{i}^{\prime}=-)$ for all $i$.
  • $w_{n}=w^{\prime}_{n}=$- and $(w_{i}=+\Rightarrow w_{i}^{\prime}=+)$ for all $i$.
It is immediate to check that this makes $\preccurlyeq_{b}$ a total order on the set of binary words of length $n$ and with alternation number at most two.
3.2. Crossing on grid graphs
In this section, we establish an elementary fact about the crossing of “monotone” curves on a grid graph. The statement and the proof are completely independent of the rest of the paper.
Consider an $(n+1)\times(n+1)$ grid digraph $D_{n}$ with “monotone” diagonals, where the vertices are the pairs $(i,j)\in[n+1]\times[n+1]$, and where there is an arc from a vertex $(i,j)$ to a distinct vertex $(i^{\prime},j^{\prime})$ when $i\leqslant i^{\prime}\leqslant i+1$ and $j\leqslant j^{\prime}\leqslant j+1$. An arc is horizontal if $j^{\prime}=j$, it is vertical if $i^{\prime}=i$, and it is diagonal otherwise. In a natural embedding, this terminology is consistent.
We introduce the “toroidal” version of $D_{n}$, which we denote $\overline{D}_{n}$. It is obtained from $D_{n}$ by identifying the vertices $(1,j)$ with the vertices $(n+1,j)$ and the vertices $(i,1)$ with the vertices $(i,n+1)$. Note in particular that the four pairs $(1,1)$, $(1,n+1)$, $(n+1,1)$, $(n+1,n+1)$ form a single vertex. With this identification, the graph $\overline{D}_{n}$ can actually be seen as embedded on a torus.
Given a directed walk and a directed cycle that are arc-disjoint, they cross whenever two consecutive arcs of the walk are separated at their common vertex by the cycle. On Figure 2, there is one crossing between the (blue) directed walk and the (red) directed cycle at column $i=9$ and row $j=4$. Crossings at vertices where the directed walk passes more than once are also possible.
We remind the reader that a walk in a directed graph is a sequence of arcs such that the head of an arc is the tail of the next arc. Moreover, a closed walk has the same definition, where the sequence is understood as a cyclic sequence.
Proposition 3.2.
Consider a closed directed walk $W$ of $\overline{D}_{n}$ that has at most $2n$ horizontal arcs and exactly $n$ vertical arcs. Then there exists a directed cycle $\Gamma$ of $\overline{D}_{n}$ using only diagonal arcs such that $W$ crosses $\Gamma$ at most once.
Proposition 3.2 is illustrated on Figure 2. Note that actually any directed cycle like $\Gamma$ in the statement, namely formed only by diagonal arcs, necessarily uses exactly $n$ arcs.
Proof of Proposition 3.2.
If $W$ has no horizontal arc, the result is immediate. Suppose thus that $W$ has at least one horizontal arc. Let’s call a directed cycle using only diagonal arcs a diagonal cycle. There are exactly $n$ diagonal cycles. Since $W$ has at least one horizontal arc and at least one vertical arc, it has in particular a vertical arc followed by a horizontal arc. These two arcs cannot be involved together in any crossing. The number of vertical arcs of $W$ being $n$, one diagonal cycle at least does not cross with $W$ at the head of a vertical arc. Call it $\Gamma$. We check that it satisfies the desired property, namely that it has at most one crossing with $W$.
Consider the universal cover of $\overline{D}_{n}$. The lifted version of $\Gamma$ is formed by copies of an infinite directed path bounding diagonal strip regions. The lifted version of $W$ is formed by copies of an infinite directed path. Pick an arbitrary copy, and keep from it a finite directed path of same length as $W$, which we denote by $\overline{W}$. We assume that $\overline{W}$ starts at a vertex at the interior of a diagonal region. Each crossing of $W$ with $\Gamma$ corresponds to $\overline{W}$ crossing the boundary of a diagonal strip region. By the choice of $\Gamma$, the directed path $\overline{W}$ can only cross such a boundary from the left to the right (i.e., by increasing the first coordinate). Since the number of horizontal arcs is at most $2n$ and the number of vertical arcs is exactly $n$, the directed path $\overline{W}$ ends either in the region where it starts, or in the region immediately on the right. Therefore, $\overline{W}$ crosses at most one boundary, which means that $W$ crosses $\Gamma$ at most once. ∎
Figure 2. Illustration of Proposition 3.2.
3.3. Theorem 3.1 as the existence of a “diagonal”
As in the statement of Theorem 3.1, consider $n$ binary words $w^{1},w^{2},\ldots,w^{n}$ of length $n$ and assume that each of them has alternation number at most two. Without loss of generality, we assume that they are indexed increasingly according to $\preccurlyeq_{b}$ (the total order defined in Section 3.1). The next result ensures the existence of a binary word as promised by Theorem 3.1, with an extra “diagonal” structure. When larger than $n$, the index $i+k$ is identified with i+kni+k-n.
Proposition 3.3.
There exists a non-negative integer $k\leqslant n-1$ such that the binary word $w_{1}^{1+k},w_{2}^{2+k},\ldots,w_{n}^{n+k}$ has alternation number at most two.
Proof.
We assume that there are at least two distinct $w^{j}$ since otherwise there is nothing to prove.
Consider the grid digraph $D_{n}$ of Section 3.2. Mark a horizontal arc $\bigl{(}(i,j),(i+1,j)\bigl{)}$ whenever $w_{i}^{j-1}\neq w_{i}^{j}$ for $j\in\{2,\ldots,n+1\}$ (with $n+1\coloneqq 1$). Mark a vertical arc $\bigl{(}(i,j),(i,j+1)\bigl{)}$ whenever $w_{i-1}^{j}\neq w_{i}^{j}$ for $i\in\{2,\ldots,n\}$. The marked arcs form the “boundary” between the symbols ++ and - in the $n\times n$ matrix where the $(j,i)$ entry is $w_{i}^{j}$; see Figure 3.
We are going to mark extra vertical and horizontal arcs. For each $j$ such that $w^{j}$ uses only one symbol, mark the vertical arc $\bigl{(}(1,j),(1,j+1)\bigl{)}$. Note that now there is exactly one vertical arc marked on each “row,” i.e., for each $j\in[n]$. The marked arcs form a subgraph $H$ of $\overline{D}_{n}$ where each vertex has equal in- and outdegree, except in two cases depicted hereafter. An all-- column (resp. all-++ column) is an index $i$ such that $w_{i}^{j}=$- (resp. $w_{i}^{j}=+$) for all $j$. There are only two cases where some vertices of $H$ may have different in- and outdegrees:
  1. (a)
    There is at least one all-- column.
  2. (b)
    There is no all-- column, but there is at least one all-++ column.
Assume that there is at least one vertex of $H$ with different in- and outdegrees. In cases (a) and (b), the definition of the order makes that $H$ has a vertex of the form $(i,j)$ that has indegree $2$ and outdegree 0, and a vertex of the form $(i^{\prime},j)$ that has indegree 0 and outdegree $2$. In case (a), $j$ is actually necessarily equal to $1$. In case (a), mark “twice” the horizontal arcs from $(i,1)$ to $(i^{\prime},1)$, i.e., the arcs
$\bigl{(}(i,1),(i+1,1)\bigr{)},\bigl{(}(i+1,1),(i+2,1)\bigr{)},\ldots,\bigl{(}(i^{\prime}-1,1),(i^{\prime},1)\bigr{)}\,$,
where, as usual, the horizontal index is counted “modulo nn.” In case (b), mark “twice” the horizontal arcs from $(i,j)$ to $(i^{\prime},j)$, i.e., the arcs
$\bigl{(}(i,j),(i+1,j)\bigr{)},\bigl{(}(i+1,j),(i+2,j)\bigr{)},\ldots,\bigl{(}(i^{\prime}-1,j),(i^{\prime},j)\bigr{)}\,.$
(In case (b), we necessarily have $i<i^{\prime}$, with the convention that $i=n+1$ means $i=1$.) We add these arcs to $H$ (each arc marked “twice” being added twice), and $H$ has now equal in- and outdegree everywhere.
By construction, every column has exactly two marked horizontal arcs (counting twice the arcs marked “twice”) and every row has exactly one vertical arc. Thus $H$ has $2n$ horizontal arcs and $n$ vertical arcs, and it is weakly connected (there is always a directed path in $H$ from the head of a vertical arc to the tail of the vertical arc in the next row). The arcs of $H$ form then a closed directed walk $W$ as in the statement of Proposition 3.2. The directed cycle $\Gamma$ whose existence is ensured by this proposition translates into the desired “diagonal” binary word: this results from an easy case checking, keeping in mind that we only have to check that a sign change translates into a crossing between $W$ and $\Gamma$, and not necessarily the other way around since we are looking for an upper bound. ∎
Figure 3. Illustration of the construction of the proof of Proposition 3.3. Each row is formed by a word of length $9$. These words are ordered from the top to the bottom according to the $\preccurlyeq_{b}$ order.